x varies directly as the square of s and inversely as t. How does x change when s is doubled? When both s and t are doubled?
Answer by Earlsdon(6294)

You can write this variation as: x = ks^2/t (k is the proportionality constant)
Now if you double s (2s), you'll have:
x = k(2s)^2/t = 4ks^2/t or exactly 4 times as much.
If you double both s (2s) and t (2t), you'll have:
x = k(2s)^2/2t = 4ks^2/2t = 2ks^2/t or exactly twice as much.
So, when s is doubled, x is quadrupled.
and when both s and t are doubled, x is doubled.